OpenStudy (anonymous):

simplify without using a calculator cos^{2}30-sin^{2}30

OpenStudy (anonymous):

So you know what cos and sin 30 are?

OpenStudy (anonymous):

sin(30)=1/2 and \[\cos =\frac{ \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

K, where u having the problem?

OpenStudy (anonymous):

the ^2

Parth (parthkohli):

Also note these two:\[\sin^2 (x) = (\sin x)^2 \\ \cos^2 (x) = (\cos x)^2\]

OpenStudy (anonymous):

Multiply by itself?

OpenStudy (anonymous):

ooooooh ok thanx

OpenStudy (anonymous):

ur welcome

Parth (parthkohli):

For example,\[\sin^2(\pi) = (\sin \pi)^2 = \cdots\]

Parth (parthkohli):

You're welcome... maybe

OpenStudy (anonymous):

thanx now i know